3. Find an interval on which Rolle’s theorem applies to f(x) = x 3 − 7x 2 + 10x. Then, find all the critical points in that interval.


1
Expert's answer
2021-04-20T07:48:07-0400

"f(x) = x^ 3 \u2212 7x^ 2 + 10x"

"=x(x^2-7x+10)=x(x-2)(x-5)"


The interval for which The rolles theorem is applicable will be [0,5].


The critical point in that interval can be calculated by putting "f'(x)=0"


"3x^2-14x+10=0\\"

"x=\dfrac{14\pm \sqrt{196-120}}{6}=\dfrac{14\pm \sqrt{76}}{6}"


Hence the critical points are "\dfrac{14- \sqrt{76}}{6},\dfrac{14+ \sqrt{76}}{6}"